OpenStudy (anonymous):

find the nth pattern of 3, 9, 15, 21, 27, ...

OpenStudy (anonymous):

this is an arithmetic series with common difference 6 nth term = a + (n-1)d = 3 + ( n - 1) 6 = 6n - 3

OpenStudy (anonymous):

is that what you require kiko? - i wasn't sure what you mean by nth pattern

OpenStudy (anonymous):

an = 3(2n-1) an = 3(4n)2 an = 9n - 6 an = 12n - 9 ur answer is correct but which of this one is correct? this are answers for that question

OpenStudy (anonymous):

the first one is the same as mine - if you distribute the 3 you get 6n - 3

OpenStudy (anonymous):

ohh mybad i didnt think it like that thanks bro

OpenStudy (anonymous):

yw